# Private Federated Learning using Preference-Optimized Synthetic Data

### v1

@@ Abstract
Machine Learning, ICML
@@
Figure 1: Left: Private Evolution (PE)-based techniques. Clients generate low-dimensional statistics which summarize the similarity of the synthetic data to their private samples. These are privately aggregated to refine the synthetic data generation for future iterations. Traditional PE (brown) uses a prompt-based method. POPri (blue) improves a naive fine-tuning method (PE+SFT, purple) by fine-tuning the LLM using preference optimization rather than fine-tuning directly on aggregated client feedback. Right: Next-token prediction accuracy on the bioRxiv dataset at privacy level $\epsilon=1$. POPri closes the accuracy gap between the fully-private and non-private settings by 68%, compared to 52% for prior synthetic data methods, and 10% for DP federated learning methods.
@@ 1 Introduction
PrE-Text (Hou et al., 2024) instead uses Private Evolution (PE) (Lin et al., 2023; Xie et al., 2024) to learn prompts that are relevant to client data. PE iteratively sends synthetic data samples to clients for feedback; each client privately measures the closeness of synthetic samples to their own data, discarding irrelevant samples. It returns this feedback to the central server, which crafts a new prompt based on the most relevant synthetic samples. Finally, an LLM uses the generated synthetic data to fine-tune a downstream model. This method of utilizing LLMs for on-device learning has some shortcomings: (1) it relies entirely on prompting to teach the LLM to generate relevant synthetic data, which may not be as effective as fine-tuning the weights. (2) It discards irrelevant samples, which may themselves contain valuable information, as shown in reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022).
In this paper, we demonstrate how to better utilize LLMs for on-device learning: we propose POPri (Preference Optimization for Private Client Data), an algorithm that reformulates synthetic data-based approaches for private on-device learning as an LLM preference optimization problem. In POPri, we directly fine-tune an LLM’s weights to improve the (DP-noised) similarity scores between generated synthetic data and private client samples. The fine-tuned LLM is used to generate synthetic data, which is used to train an on-device model.
@@
(1) We propose POPri, a novel method that casts private on-device learning under the synthetic data framework as an LLM preference optimization problem. Prior work in this space relied on PE, which uses client feedback exclusively to generate new prompts (Hou et al., 2024; Xie et al., 2024). We alter this feedback to instead provide client preferences, and subsequently exploit recent advances in preference optimization (Rafailov et al., 2023). This recasting allows us to more effectively exploit the capabilities of LLMs for on-device learning problems.
@@
(3) We demonstrate the utility of POPri on this new benchmark set of datasets, as well as a dataset collected from PubMed (Yu et al., 2023; Xie et al., 2024). Across all datasets, POPri achieves the best downstream metrics. For example, Figure 1 shows that on our bioRxiv dataset at a privacy level of $\epsilon=1.0$, POPri outperforms PE-based algorithms by 2 full percentage points, and closes the gap between fully private and non-private baselines by over 68%, compared to 52% for PE. It outperforms DP-FL-based methods by even more. Additional experimental details, results, and ablations are provided in Section 5.
@@ 2.1 Problem Statement
User-level (distributed) differential privacy (DP). We say two datasets $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are neighboring if they differ in at most one client’s data. That is, there exists an $i\in[n]$ such that for all $j\neq i$, $S_{j}=S^{\prime}_{j}$. A randomized mechanism $\mathcal{M}$ is ($\epsilon$, $\delta$)-DP if, for any pair of neighboring datasets $\mathcal{S}$, $\mathcal{S^{{}^{\prime}}}$ that differ by one sample and any possible output set $E$, it holds that ${\rm Pr}[\mathcal{M}(\mathcal{S})\in E]\leq e^{\epsilon}{\rm Pr}[\mathcal{M}(\mathcal{S^{{}^{\prime}}})\in E]+\delta$. The post-processing property of a DP mechanism ensures that any data-independent transformation applied to its output preserves the same DP guarantees. (Dwork, 2006; Dwork & Roth, 2014).
@@
The server seeks an algorithm to optimize the downstream next-word prediction performance of $\tilde{\Phi}$ on a test set of private client data, subject to an $(\epsilon,\delta)$-DP constraint.
@@ 2.2 Related Work
One approach for generating synthetic text data is to fine-tune an LLM (with DP-SGD) on private data (Kurakin et al., 2023; Yu et al., 2024) and then using the LLM to generate synthetic data. However, client hardware constraints render this approach infeasible on-device. Recent works have relied instead on privacy-aware prompt engineering to generate synthetic data (Wu et al., 2024; Xie et al., 2018; Hou et al., 2024). An important framework by Lin et al. (2023) called Private Evolution (PE) is the basis for several competitive DP synthetic text algorithms, including Aug-PE (Xie et al., 2024) and PrE-Text (Hou et al., 2024). Roughly, these algorithms use the public LLM $\Psi$ to generate synthetic data, score each synthetic data according to its closeness to the client data, and discard synthetic data with low scores. The surviving synthetic data are used as in-context examples for $\Psi$ to generate synthetic data. Private Evolution may sacrifice data quality in two ways: First, it uses in-context learning, which is often less effective than fine-tuning (Mosbach et al., 2023). Second, discarding low-score synthetic data may lose useful information (Ouyang et al., 2022). We address both by turning the DP synthetic generation problem into an LLM preference optimization problem.
@@ 3 POPri
The core idea of POPri (Preference Optimization for Private Client Data) is a natural reformulation of private on-device learning from synthetic data as an LLM preference optimization problem, which enables the use of powerful LLM alignment methods like DPO (Rafailov et al., 2023). In this section, we detail the POPri design principles and algorithm. POPri’s design is based on two related questions.
@@ 2. How should we use client feedback?
Figure 2: 2-PCA visualization of synthetic data from POPri and PE+SFT, and evaluation data. We see that POPri’s synthetic data distribution (left) is much closer to the evaluation data distribution (right) than the PE+SFT synthetic data distribution (middle). Naively fine-tuning with SFT on PE-generated synthetic data does not make best use of client feedback.
@@
(3) Preference Optimization (PO). Preference optimization-based methods like DPO (Rafailov et al., 2023) instead directly optimize the LLM to produce higher-scoring samples (where the score can be defined by the user of the algorithm) using preference pairs. In other words, they are designed to directly make use of the low-dimensional scores we collect from client feedback. Hence, we expect such methods to produce higher quality synthetic data, as evaluated on downstream tasks. Other choices for similar methods included RLHF (Ouyang et al., 2022), which we avoided due to the high computational requirements of training a reward model, and IPO (Gheshlaghi Azar et al., 2024) which we did not choose as a result of an ablation in Section E.4.
@@
1:  Input: Clients private data $\{S_{i}\}_{i\in[n]}$, Number of rounds $T$, Number of generated samples $N_{\rm syn}$, Noise multiplier $\sigma$, LLM $\Psi$, embedding model $\Gamma$, base prompt $\eta$, participating clients in each round $\mathcal{S}^{t}$, “rejected” index $\ell$, initial sample set $\Omega$, number of clients sampled $L$
@@
10:      Generate prompt $\eta_{k}=\Psi(\eta,\Omega)$,
@@
20:    Secure aggregate client scores: ${\rm Scores}_{t}=\frac{1}{n}\sum_{i}{\rm Scores}_{i,t}$
@@ 3.1 POPri Algorithm
1. Initial sample population.
We start with an initial set of samples $\Omega$, which come from a publicly available source. This source may be text that is publicly available on the internet or text generated by a publicly available LLM.
@@
We create $K$ prompts. A prompt is generated by randomly sampling three samples from $\Omega$ and prompting LLaMA-3-8B (Touvron et al., 2023) to generate a fourth sample given the first three samples as examples. The exact prompt is given in Appendix B. For each of the $K$ prompts, we generate $J$ synthetic samples (by running the prompt independently $J$ times). In total, the server generates $K\times J$ synthetic samples and sends them to every client in $\mathcal{S}^{t}$, i.e., the clients sampled in round $t$.
@@
Next, each client in $\mathcal{S}^{t}$ scores the synthetic samples. Specifically, each client calculates the average cosine similarity between each of the $K\times J$ synthetic samples and the entire client dataset (Algorithm 2). The use of cosine similarity differs from PE, which uses a nearest neighbors histogram (Lin et al., 2023; Hou et al., 2024; Xie et al., 2024)–using cosine similarity is critical to the performance of POPri as we found in our ablations (see Section 5.2). These similarities for every synthetic sample are arranged into a vector. We clip this vector to a norm of 1, which caps the contribution of each client (similar to how gradient updates are clipped per client in DP-FL (McMahan et al., 2017b)). Clipping is done primarily for privacy reasons, as we will elaborate later. Clipping also ensures that the contribution of clients with large amounts of data does not overwhelm the contribution of clients with small amounts of data. We then add $\mathcal{N}(0,\sigma^{2}I/L)$ (where $I$ is the identity matrix of size $KJ\times KJ$) noise to the resulting vector to ensure DP ($\sigma^{2}$ controls the $(\epsilon,\delta)$). Finally, we aggregate scores via secure aggregation (Bonawitz et al., 2016), yielding a DP score for each synthetic sample that reflects its relevance to client data.
[…]
@@
bioRxiv DP-FedAvg Original DistilGPT2 72.2 61.7 61.7 60.6
DP-FTRL Original 61.8 61.8
PE Synthetic 66.2 66.3
PE + SFT Synthetic 64.8 64.6
POPri (ours) Synthetic 68.6 68.6
Congress DP-FedAvg Original DistilGPT2 74.5 69.2 69.2 68.4
DP-FTRL Original 69.1 69.1
PE Synthetic 70.3 70.4
PE + SFT Synthetic 70.0 70.2
POPri (ours) Synthetic 71.3 71.3
Dataset Method Data Type On-device Model $\epsilon=\infty$ $\epsilon=4$ $\epsilon=1$
PubMed (Yue et al., 2023b) PE GPT-2-Large, Synthetic (2000) ${\textsc{BERT}_{\rm small}}$ 47.6 27.9 27.2
PE Llama-2-7b-chat-hf, Synthetic (2000) 27.5
@@
POPri (ours) Synthetic (2000) 29.2 29.4
4. LLM Preference Optimization.
The key insight of our paper is that by generating $J$ synthetic samples from $K$ prompts and scoring all of them using DP client feedback, we can create a preference dataset where for each of the $K$ prompts, we can assemble a “good sample” and a “bad sample”. This design choice allows the usage of powerful LLM preference optimization algorithms (we choose DPO (Rafailov et al., 2023)) to finetune the LLM $\Psi$. In detail, each of the $K$ prompts have $J$ synthetic samples which are ranked according to the scores we gathered. Then for each of the $K$ prompts, we set the highest scoring sample as the “chosen sample” and the $\ell$-th highest scoring sample as the “rejected sample”. This resulting preference dataset can then be passed, along with the LLM $\Psi$, into the DPO preference optimization loss (Rafailov et al., 2023):
@@ 4 LargeFedBench: A Federated Benchmark for LLM Evaluation
Figure 3: Next-token prediction accuracy performance of four methods as a function of the fraction of clients used per round of training. POPri improves performance and decreases sensitivity to the fraction of clients used in each round. The left panels have a different number of clients per round with the same total number of clients, and the right panels have a different total number of clients with the same number of clients per round. The top and bottom use different datasets and all methods have a privacy budget $(\epsilon,\delta)$-DP = (1, $3\times 10^{-6}$).
@@ 5 Experiments
Datasets.
We evaluate POPri on the LargeFedBench datasets (Congress and bioRxiv), as well as the PubMed (Yu et al., 2023; Xie et al., 2024) used in the evaluation of Private Evolution (Aug-PE) (Xie et al., 2024). PubMed contains abstracts of medical papers published between August 1-7, 2023 (details in Section D.2.2).
@@
For the LLM $\Psi$, we use LLaMA-3-8B, which has a knowledge cutoff date of March 2023 (AI@Meta, 2024). For embedding models (used in measuring semantic distance between text samples), we use ‘all-MiniLM-L6-v2’ sentence transformer (Reimers & Gurevych, 2019). We choose DistilGPT2 (Sanh et al., 2019) as the downstream on-device language model, which has only 82M parameters. For synthetic text generation (using the LLM $\Psi$), we set the maximum sequence length to 64 for the bioRxiv and Congressional Speeches evaluations and 512 for PubMed, unless otherwise specified. During training, we evaluate the models on the validation dataset and select the checkpoint that achieves the best validation performance as the model that is evaluated on the test set.
@@
We primarily evaluate each method under next-token prediction accuracy of the final downstream on-device model $\tilde{\Phi}$. In some ablations we also measure the distance of the synthetic dataset to the private dataset using the Fréchet Inception Distance (FID) (Heusel et al., 2017).
@@
We compare POPri with several baselines: (1) DP-FedAvg (McMahan et al., 2017a) (2) DP-FTRL (Kairouz et al., 2021a) (3) Private Evolution (PrE-Text (Hou et al., 2024) and Aug-PE (Xie et al., 2024)). DP-FedAvg and DP-FTRL directly privately fine-tune the downstream model $\Phi$ on the client data. Private Evolution (PrE-Text and Aug-PE) generates synthetic data on which the downstream on-device model $\Phi$ is finetuned. Note that on the PubMed dataset, we compare to Aug-PE results obtained with models of similar size to the model we use (around 8B parameters). [^3] This comparison is somewhat unfair, as the models they use are older. They obtain better results using GPT-3.5, a 175B parameter LLM, 20$\times$ larger than the one we use. We also include $\epsilon=0$ (fully private) and $\epsilon=\infty$ (fully non-private) baselines. The $\epsilon=0$ baseline evaluates the public DistilGPT2 checkpoint on the test sets with no further fine-tuning. The $\epsilon=\infty$ baseline is DistilGPT2 fine-tuned on the private training set directly and evaluated directly on the private test set. More details about the setup can be found in Appendices C and D.2.
Figure 4: PCA visualization of POPri synthetic data embeddings over rounds. Right (6) Panels: PCA-2 plots for synthetic data and evaluation data embeddings from the best checkpoint each round for 20 iterations. The orange (round 7) and maroon point clouds represent the round with the lowest FID score and the validation dataset, respectively. Top Left Panel: FID score vs. rounds. Bottom Left Panel: Median distance to the medoid vs rounds. Running POPri for too many rounds appears to cause overfitting.
@@
Reduction factor (FedAvg / POPri) 11.71×11.71$\times$11.71 × 4555×4555$\times$4555 × 480.0×480.0$\times$480.0 ×
Reduction factor (PE / POPri) $0$ $0$ 0.270×0.270$\times$0.270 × 0.024×0.024$\times$0.024 ×
Table 2: Table setting. Communication and compute cost comparison per round (and per-client for download/upload/client runtime cost) across methods on the bioRxiv dataset with 1000 clients sampled per round. Download and upload are measured in floats; runtimes are measured in GPU seconds (lower is better). “Reduction factor ($X$ / POPri)” is the cost of method $X$ divided by the cost of POPri for the given resource. Server runtime for FedAvg is left blank as it is negligible compared to other methods. Overall, we view POPri as suitable when server compute is relatively cheap, and improved sample quality is important enough to justify higher on-device communication and computation costs relative to PE (Table 1).
@@
All baselines use a privacy guarantee of ($\epsilon$, $\delta$)-DP where $\delta$=3×$\times 10^{-6}$ and $\epsilon$=1 or $\epsilon$=7 for each of the bioRxiv and Congressional Speeches datasets. For PubMed, we set $\delta={1/[N_{priv}\cdot{\rm log}(N_{priv})]}$ ($N_{priv}$ is the number of private samples) and $\epsilon$=1 or $\epsilon$=4 and fine-tune ${\rm BERT}_{small}$ for fair comparison to the results from Xie et al. (2024). We follow the privacy accounting method detailed in Section 3 for POPri. Details for all baselines are in Section D.1.

### v2

# POPri: Private Federated Learning using Preference-Optimized Synthetic Data

@@ Abstract
differential privacy, synthetic data, federated learning, preference optimization
@@
Figure 1: Left: Private Evolution (PE)-based techniques. Clients generate low-dimensional statistics which summarize the similarity of the synthetic data to their private samples. These are privately aggregated to refine the synthetic data generation for future iterations. Traditional PE (brown) uses a prompt-based method. POPri (blue) improves a naive fine-tuning method (PE+SFT, purple) by fine-tuning the LLM using policy optimization rather than fine-tuning directly on aggregated client feedback. Right: Next-token prediction accuracy on the bioRxiv dataset at privacy level $\epsilon=1$. POPri closes the accuracy gap between the fully-private and non-private settings by 58%, compared to 23% for prior synthetic data methods, and 3% for DP federated learning methods.
@@ 1 Introduction
PrE-Text (Hou et al., 2024) instead uses Private Evolution (PE) (Lin et al., 2023; Xie et al., 2024; Lin et al., 2025) to learn prompts that are relevant to client data. PE iteratively sends synthetic data samples to clients for feedback; each client privately measures the closeness of synthetic samples to their own data, discarding irrelevant samples. It returns this feedback to the central server, which crafts a new prompt based on the most relevant synthetic samples. Finally, an LLM uses the generated synthetic data to fine-tune a downstream model. This method of utilizing LLMs for on-device learning has some shortcomings: (1) it relies entirely on prompting to teach the LLM to generate relevant synthetic data, which may not be as effective as fine-tuning the weights. (2) It discards irrelevant samples, which may themselves contain valuable information, as shown in reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022).
In this paper, we demonstrate how to better utilize LLMs for on-device learning: we propose POPri (Policy Optimization for Private Data), an algorithm that reformulates synthetic data-based approaches for private learning as an LLM policy optimization problem. In POPri, we directly fine-tune an LLM’s weights to improve the (DP-noised) similarity scores between generated synthetic data and private client samples. The fine-tuned LLM is used to generate synthetic data, which is used to train a downstream model.
@@
(1) We propose POPri, a novel method that casts private learning under the synthetic data framework as an LLM policy optimization problem. Prior work in this space relied on PE, which uses client feedback exclusively to generate new prompts (Hou et al., 2024; Xie et al., 2024). We alter this feedback to instead provide client rewards, and subsequently exploit recent advances in policy optimization (Rafailov et al., 2023). This recasting allows us to more effectively exploit the capabilities of LLMs for on-device learning problems.
@@
(3) We demonstrate the utility of POPri on this new benchmark set of datasets, as well as two central (i.e., the setting where all data is present on the server and no server-client communication is needed) DP benchmarks from prior work (Yu et al., 2023; Xie et al., 2024). Across all datasets and tasks (we consider next token prediction and text classification), POPri achieves the best downstream metrics. For example, Figure 1 shows that on our bioRxiv dataset at a privacy level of $\epsilon=1.0$, POPri outperforms PE-based algorithms by 6 full percentage points, and closes the gap between fully private and non-private baselines by over 58%, compared to 23% for PE. It outperforms DP-FL-based methods by even more. Additional experimental details, results, and ablations are provided in Section 5.
@@ 2.1 Problem Statement
User-level (distributed) differential privacy (DP). We say two datasets $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are neighboring if they differ in at most one client’s data. That is, there exists an $i\in[n]$ such that for all $j\neq i$, $S_{j}=S^{\prime}_{j}$. A randomized mechanism $\mathcal{M}$ is ($\epsilon$, $\delta$)-DP if, for any pair of neighboring datasets $\mathcal{S}$, $\mathcal{S^{{}^{\prime}}}$ that differ by an entire client’s data and any possible output set $E$, it holds that ${\rm Pr}[\mathcal{M}(\mathcal{S})\in E]\leq e^{\epsilon}{\rm Pr}[\mathcal{M}(\mathcal{S^{{}^{\prime}}})\in E]+\delta$. The post-processing property of a DP mechanism ensures that any data-independent transformation applied to its output preserves the same DP guarantees. (Dwork, 2006; Dwork & Roth, 2014).
We also evaluate on central DP baselines, so we define central DP below; in that case the final model $\tilde{\Phi}$ should protect central DP:
Central (example-level) differential privacy (DP). We say two datasets (both fully present on the server) $S=\{s_{1},\ldots,s_{m}\}$ and $S^{\prime}=\{s_{1}^{\prime},\ldots,s_{m}^{\prime}\}$ are neighboring if they differ in at most one example’s data. That is, there exists an $i\in[m]$ such that for all $j\neq i$, $s_{j}=s^{\prime}_{j}$. A randomized mechanism $\mathcal{M}$ is ($\epsilon$, $\delta$)-DP if, for any pair of neighboring datasets $\mathcal{S}$, $\mathcal{S^{{}^{\prime}}}$ that differ by one sample and any possible output set $E$, it holds that ${\rm Pr}[\mathcal{M}(S)\in E]\leq e^{\epsilon}{\rm Pr}[\mathcal{M}(S^{\prime})\in E]+\delta$.
@@
The server seeks an algorithm to optimize the downstream performance (in our paper, this is either next token prediction accuracy or text classification accuracy) of $\tilde{\Phi}$ on a test set of private data, subject to an $(\epsilon,\delta)$-DP constraint.
@@ 2.2 Related Work
One approach for generating synthetic text data is to fine-tune an LLM (with DP-SGD) on private data (Kurakin et al., 2023; Yu et al., 2024) and then using the LLM to generate synthetic data. However, client hardware constraints render this approach infeasible on-device. Recent works have relied instead on privacy-aware prompt engineering to generate synthetic data (Wu et al., 2024; Xie et al., 2018; Hou et al., 2024). An important framework by Lin et al. (2023, 2025) called Private Evolution (PE) is the basis for several competitive DP synthetic text algorithms, including Aug-PE (Xie et al., 2024) and PrE-Text (Hou et al., 2024). Roughly, these algorithms use the public LLM $\Psi$ to generate synthetic data, score each synthetic data according to its closeness to the client data, and discard synthetic data with low scores. The surviving synthetic data are used as in-context examples for $\Psi$ to generate synthetic data. In concurrent work to ours, Zou et al. extend the PE framework to generate synthetic data from multiple pretrained language models (LMs), and present “good” and “bad” responses to the LMs in the next round for in-context learning (Zou et al., 2025). Private Evolution may sacrifice data quality in two ways: First, it uses in-context learning, which is often less effective than fine-tuning (Mosbach et al., 2023). Second, discarding low-score synthetic data may lose useful information (Ouyang et al., 2022). We address both by turning the DP synthetic generation problem into an LLM policy optimization problem.
@@ 3 POPri
The core idea of POPri (Policy Optimization for Private Data) is a natural reformulation of private on-device learning from synthetic data as an LLM policy optimization problem, which enables the use of powerful LLM alignment methods like DPO (Rafailov et al., 2023). In this section, we detail the POPri design principles and algorithm. POPri’s design is based on two related questions.
@@ 2. How should we use client feedback?
Figure 2: 2-PCA visualization of synthetic data from POPri and PE+SFT, and evaluation data. We see that POPri’s synthetic data distribution (left) is much closer to the evaluation data distribution (right) than the PE+SFT synthetic data distribution (middle). Naive fine-tuning with SFT on PE-generated synthetic data does not make best use of client feedback.
@@
(3) Policy Optimization (PO). Policy optimization-based methods like DPO (Rafailov et al., 2023) instead directly optimize the LLM to produce higher-scoring samples (where the score can be defined by the user of the algorithm). In other words, they are designed to directly make use of the low-dimensional scores we collect from client feedback. Hence, we expect such methods to produce higher quality synthetic data, as evaluated on downstream tasks.
@@
1:  Input: Clients private data $\{S_{i}\}_{i\in[n]}$, Number of rounds $T$, Number of generated samples $N_{\rm syn}$, Noise multiplier $\sigma$, LLM $\Psi$, embedding model $\Gamma$, base prompt $\eta$, participating clients in each round $\mathcal{S}^{t}$, “rejected” index $\ell$, random prompt generator $\Lambda(\cdot)$, number of clients sampled $L$
@@
10:   Generate prompt $\eta_{k}=\Lambda(\eta)$,
@@
20:  Secure aggregate scores: ${\rm Scores}_{t}=\frac{1}{L}\sum_{i\in\mathcal{S}^{t}}{\rm Scores}_{i,t}$
@@ 3.1 POPri Algorithm
We generate $K$ prompts (details in Section B.2). A prompt is generated by randomly sampling three samples from $\Omega$ and prompting LLaMA-3-8B (Touvron et al., 2023) to generate a fourth sample given the first three samples as examples. The exact prompt is given in Appendix B. For each of the $K$ prompts, we generate $J$ synthetic samples (by running the prompt independently $J$ times). In total, the server generates $K\times J$ synthetic samples, embeds them using a small sentence embedding model $\Gamma$ and sends the embeddings to every client in $\mathcal{S}^{t}$, i.e., the clients sampled in round $t$.
@@
Next, each client in $\mathcal{S}^{t}$ scores the synthetic samples. Specifically, each client calculates, for each of the $K\times J$ synthetic samples, its cosine similarity with each of the client’s private samples, averaged over the client’s samples (Algorithm 4). The use of cosine similarity differs from PE, which uses a nearest neighbors histogram (Lin et al., 2023; Hou et al., 2024; Xie et al., 2024)–using cosine similarity is critical to the performance of POPri as we found in our ablations (see Section 5.2). These similarities for every synthetic sample are arranged into a vector. We clip this vector to a norm of 1, which caps the contribution of each client (similar to how gradient updates are clipped per client in DP-FL (McMahan et al., 2017b)). Clipping is done primarily for privacy reasons, as we will elaborate later. Clipping also ensures that the contribution of clients with large amounts of data does not overwhelm the contribution of clients with small amounts of data. We then add $\mathcal{N}(0,\sigma^{2}I/L)$ (where $I$ is the identity matrix of size $KJ\times KJ$) noise to the resulting vector to ensure DP ($\sigma^{2}$ controls the $(\epsilon,\delta)$). Finally, we aggregate scores via secure aggregation (Bonawitz et al., 2016), yielding a DP score for each synthetic sample that reflects its relevance to client data.
Table 1: Accuracy ($\%,\uparrow$) of different algorithms on a variety of tasks and datasets (bioRxiv, Congress, PubMed are next-token-prediction accuracy, OpenReview is text classification accuracy). The highest accuracy across all methods is in $\bf bold$}. All standard deviation error bars are less than 0.5.
@@
bioRxiv DP-FedAvg Original DistilGPT2 41.5 29.0 28.3 27.9
DP-FTRL Original 29.0 28.2
PE Synthetic 31.0 31.1
PE + SFT Synthetic 28.6 28.6
POPri (ours) Synthetic 34.4 34.8
Congress DP-FedAvg Original DistilGPT2 35.7 29.1 29.0 26.9
DP-FTRL Original 29.1 29.0
PE Synthetic 27.3 27.0
PE + SFT Synthetic 27.1 27.1
POPri (ours) Synthetic 30.6 30.4
PubMed (Yue et al., 2023b) PE Llama-2-7b-chat-hf, Synthetic (2000) $\textsc{BERT}_{\rm small}$} 47.6 27.5
@@
POPri (ours) Synthetic (2000) 29.4
OpenReview (Xie et al., 2024) PE Llama-2-7b-chat-hf, Synthetic (2000) $\textsc{RoBERTa}_{\rm base}$} 50.8 37.0 32.0
PE Opt-6.7b, Synthetic (2000) 32.1
POPri (ours) Synthetic (2000) 40.2
3. LLM Policy Optimization.
The key insight of our paper is that by generating $J$ synthetic samples from $K$ prompts and scoring all of them using DP client feedback, we can create a preference dataset where for each of the $K$ prompts, we can assemble a “good sample” and a “bad sample”. This design choice allows the usage of powerful LLM policy optimization algorithms (we choose DPO (Rafailov et al., 2023)) to finetune the LLM $\Psi$. In detail, each of the $K$ prompts have $J$ synthetic samples which are ranked according to the scores we gathered. Then for each of the $K$ prompts, we set the highest scoring sample as the “chosen sample” and the $\ell$-th highest scoring sample as the “rejected sample”. This resulting preference dataset can then be passed, along with the LLM $\Psi$, into the DPO preference optimization loss (Rafailov et al., 2023):
@@ 4 LargeFedBench: A Federated Benchmark for LLM Evaluation
Figure 3: Next-token prediction accuracy performance of four methods as a function of the number of clients sampled per round out of 10000. We see that across different client participation scenarios, POPri consistently performs the best.
@@ 5 Experiments
Datasets and tasks.
For next token prediction accuracy, we evaluate POPri on the LargeFedBench datasets (Congress and bioRxiv), as well as PubMed (Yu et al., 2023; Xie et al., 2024) used in the evaluation of Private Evolution (Aug-PE) (Xie et al., 2024). PubMed contains abstracts of medical papers published between August 1-7, 2023 (details in Section D.2.2). For text classification, we evaluate POPri on OpenReview consisting of ICLR 2023 reviews published on November 5, 2022 which was used in the evaluation of Private Evolution (Aug-PE) (Xie et al., 2024). Note that PubMed and OpenReview are evaluations in the central DP setting, where the entire dataset is present on the server and no server-client communication is needed. To execute POPri on PubMed, we use the central DP version of POPri, detailed in Algorithm 2. For OpenReview, we employ conditional generation [^3] Simply speaking, we score synthetic data generated for each class separately but combine the preference datasets into one to finetune the LLM $\Psi$. (similar to PE (Lin et al., 2023; Xie et al., 2024), where the generation is conditioned on being given a class to generate. The (central) conditional generation version of POPri is detailed in Algorithm 3.
@@
Next token prediction tasks: We use LLaMA-3-8B for the LLM $\Psi$ (Grattafiori et al., 2024), which has a knowledge cutoff date of March 2023 (AI@Meta, 2024). For embedding models (used in measuring semantic distance between text samples), we use ‘all-MiniLM-L6-v2’ sentence transformer (Reimers & Gurevych, 2019b). We choose DistilGPT2 (Sanh et al., 2019) as the downstream on-device language model for the LargeFedBench evaluations, which has only 82M parameters, and $\text{BERT}_{\text{small}}$ as the downstream model for the PubMed evaluation to be consistent with Xie et al. (2024). For synthetic text generation (using the LLM $\Psi$), we set the maximum sequence length to 64 for the bioRxiv and Congressional Speeches evaluations and 512 for PubMed/OpenReview.
Text classification: We use LLaMA-2-7b-chat-hf for the LLM $\Psi$ (Touvron et al., 2023) to ensure our evaluation was not contaminated, as the knowledge cutoff for LLaMA-2-7b-chat-hf is September 2022 (before the publish date of the ICLR 2023 reviews). For the embedding model we use the ‘sentence-t5-xl’ sentence transformer (Reimers & Gurevych, 2019a), and use $\text{RoBERTa}_{\text{base}}$ as the downstream model to be consistent with Xie et al. (2024).
@@
We primarily evaluate each method on accuracy (next-token or text classification) of the final downstream on-device model $\tilde{\Phi}$. In some ablations we also measure the distance of the synthetic dataset to the private dataset using the Fréchet Inception Distance (FID) (Heusel et al., 2017). During training, we evaluate the models on the validation dataset and select the checkpoint that achieves the best validation performance as the model that is evaluated on the test set.
@@
We compare POPri to several baselines: (1) DP-FedAvg (McMahan et al., 2017a) (2) DP-FTRL (Kairouz et al., 2021a) (3) Private Evolution (PrE-Text (Hou et al., 2024) and Aug-PE (Xie et al., 2024)). DP-FedAvg and DP-FTRL directly privately fine-tune the downstream model $\Phi$ on the client data. Private Evolution (PrE-Text and Aug-PE) generates synthetic data on which the downstream on-device model $\Phi$ is finetuned. Note that on the PubMed and OpenReview dataset, we compare to Aug-PE results obtained with models of similar size to the model we use (7B-8B parameters) and which are not potentially contaminated (i.e. model was possibly trained on the benchmark dataset). We also include $\epsilon=0$ (fully private) and $\epsilon=\infty$ (fully non-private) baselines. The $\epsilon=0$ baseline for the LargeFedBench evaluations evaluates the public DistilGPT2 checkpoint on the test sets with no further fine-tuning. The $\epsilon=\infty$ baseline is the downstream model finetuned directly on the private training set centralized on the server with no noise. The $\epsilon=0$ baseline for OpenReview is the accuracy obtained by predicting everything to be the most populous class. More details about the setup can be found in Appendices C and D.2.
Figure 4: PCA visualization of POPri synthetic data embeddings over rounds. Right 6 Panels: PCA-2 plots for synthetic data and evaluation data embeddings from the best checkpoint each round for 20 iterations. The orange (round 7) and maroon point clouds represent the round with the lowest FID score and the validation dataset, respectively. Top Left Panel: FID score vs. rounds. Bottom Left Panel: Median distance to the medoid vs rounds. Running POPri for too many rounds appears to cause overfitting.
@@
Reduction factor (FedAvg / POPri) $11.71\times$ $4555\times$ $480.0\times$
Reduction factor (PE / POPri) $\times 0.100×0.100\times$ $270\times 0.0$$024\times$
Table 2: Table setting. Communication and computation cost comparison per round (and per-client for download/upload/client runtime cost) across methods on the bioRxiv dataset with 1000 clients sampled per round. Download and upload are measured in floats; runtimes are measured in GPU seconds (lower is better). “Reduction factor ($X$ / POPri)” is the cost of method $X$ divided by the cost of POPri for the given resource; green is a reduction, red is an increase. Server runtime for FedAvg is left blank as it is negligible compared to other methods. Overall, we view POPri as suitable when server compute is relatively cheap, and improved sample quality is important enough to justify higher on-device communication and computation costs relative to PE (Table 1).
@@
[…]